# The impact of assembly history on the X-ray detectability of halos

### v1

@@ 3.1 Build-up of the gas content
One might wonder how representative the hot gas fraction is of the total gas content in a group or cluster of galaxies. To address this, we analyse the expected hot gas fraction (hot gas mass over halo mass within $R_{500}$) as a function of the total gas fraction (hot and cold gas over halo mass within $R_{500}$) in Fig. 4. It is worth mentioning that Magneticum’s predictions on the $f_{\mathrm{gas}}-M_{halo}$ relation are in excellent agreement with the median and scatter of the hot gas fraction from eROSITA stacked groups (Popesso et al. 2024b) and the XXL groups and clusters with halo mass measurements from weak lensing (Akino et al. 2022). We divide the LC30 sample into logarithmic halo mass bins of $M_{500}$, showing their scatter (coloured shaded regions), and plot the 1:1 scaling relation in dashed black. If the relation follows the 1:1 line this implies that all gas in the halo is hot. In the lateral panels, we plot the number counts of halos in the bins we used to calculate the median and error.
We deduce that if a halo has a low total gas fraction compared to haloes with a similar mass, most (if not all) of it will be in the hot phase. The gravitational potential well-compresses the hot plasma, increasing its kinetic energy, which is later converted into thermal energy. As a result, larger halos tend to heat more gas. In the clusters regime, this process converts most of the gas into a hot, X-ray-bright component. However, in less massive halos, the colder (star-forming) gas component can survive longer within $R_{500}$. Consequently, a larger total gas fraction does not necessarily imply a well-mixed hot halo. We argue that measuring only the X-ray-emitting gas in the group regime ($M_{500}<10^{13.5}M_{\odot}$) could lead to a significant underestimation of the total gas fraction (up to 50%) if this effect is not considered – at least in the $f_{\mathrm{gas,\,tot}}>0.05$ regime of our simulations. This result is robustly predicted by the number of halos that are estimated to have $f_{\mathrm{gas,\,tot}}$ in the regime $0.05-0.10$, which represent $\geq 50\%$ of the groups in $M_{500}<10^{13.5}M_{\odot}$ (see top histograms). In this regard, recent studies on the gas fraction have called for stronger AGN feedback than most hydrodynamical simulations (but not Magneticum, see Popesso et al. 2024c) based on evidence collected from the cosmic shear, X-ray, and SZ measurements (e.g. Preston et al. 2023; Amon et al. 2023; Bigwood et al. 2024; McCarthy et al. 2024; Hadzhiyska et al. 2024; Ferreira et al. 2024; La Posta et al. 2024). However, we highlight here the need to account for complementary techniques to include cold gas (e.g. measuring the MgII covering factor; Anand et al. 2021) in Milky Way-mass halos.
@@
Furthermore, the mass accretion history is strikingly different in the two samples (right panels in Fig. 5). We overplot the self-similar evolution of a DM halo in the given mass range at $z=0$ expected in the range of concentration $c_{500}$ [^4] The DM concentration $c_{\Delta}$ defines how centrally concentrated the DM is in a halo. Analytically it reads $c_{\Delta}=R_{\Delta}/R_{s}$ where $R_{s}$ is the scale radius in an NFW profile (Navarro et al. 1997). of our simulated halos $103<c_{500}<$ 10 (Diemer et al. 2013; Diemer 2018). Faint halos typically gain most of their mass early on, with later phases marked by gas depletion due to star formation, AGN activity, and stellar feedback. Interestingly, the halo mass growth experiences a slow-down in the faint sample at the epoch when the baryon fraction begins to decline. Such decline can be indeed connected with a general failure in replenishing the gas content in the halo which will reduce in the future due to various feedback and star formation. However, their overall evolution is consistent with the pseudo-evolution of a stationary NFW density profile (shown in Fig. 5), changing due to the evolution of a halo mass defined by a density threshold $\Delta=500$ that decreases with time. In contrast, bright halos exhibit steady exponential accretion at all times, removing the $1$ dex difference in halo mass originally present at $z=3.5$. A visual inspection of the bright halos at this epoch reveals that they undergo major mergers, whereas the faint ones exhibit very few substructures and display a regular, symmetric X-ray emission around the central galaxy –reminiscent of fossil groups (Aguerri & Zarattini 2021).
@@
Figure 7: Cumulative density distribution of the normalised density contrast for bright (orange) and faint (blue) halos at three redshifts. The solid lines are the median of the bootstrapped sample (shaded lines) for halos within $10^{13}-10^{13.5}M_{\odot}$ selected at $z=0$. From left to right, the normalised density contrast is calculated for different radii, namely 1, 2, and 5 cMpc/h. In the top-left corner of each panel, we also report its correspondence to the median $R_{500}$ in the sample.
@@ 3.2 The role of the central SMBH
The activity of the central and satellite SMBHs can severely impact halos across their cosmic time. While AGN feedback heats and displaces the gas, altering the halo’s thermodynamic state (Eckert et al. 2021), many simulations (Steinborn et al. 2016, 2018; Castro et al. 2021) have shown it is insufficient to eject gas entirely from the halo at late times. In reality, most of the strong AGN activity episodes happen at the early phases of the halos assembly ($z=2-3$, see for example Castro et al. 2021; Ragagnin et al. 2022; Ayromlou et al. 2023) which can cause an efficient displacement of the gas well beyond the virial radius at the group scales. On the other hand, DM is less sensitive to the gas displacement (and completely insensitive to the thermal feedback), leading to a continuous accretion even at later times. Biffi et al. (2018) and Marini et al. (2024, 2025b) showed that Magneticum correctly reproduces the AGN luminosity function at $z=0$ and in LC30, respectively. Therefore, we expect the simulated AGN population to be representative of the real one.
@@
The BH evolution demonstrates that the feedback activity of the faint sample steadily declines over time, with no recent episodes accounting for the decrease of the gas content around $z\simeq 0-1.5$. In contrast, the bright sample exhibits a rise in AGN activity at the epoch when the baryonic fraction changes ($z\simeq 1-1.5$). If a major merger occurs at this redshift, the instantaneous availability of gas in the halo fuels the central SMBH, triggering a thermal energy release in the bright sample. Due to the subgrid model implemented, the BH accretion is more efficient and less turbulent during the accretion of cold gas particles, although both cold and hot gas phases contribute to the feeding process. Because cold gas is accreted more efficiently, (i) it is quickly depleted and (ii) it raises the temperature significantly more than hot gas alone. In response, the SMBH releases thermal energy, which, while insufficient to fully deplete the gas reservoir of the halo, significantly increases the hot gas fraction.
@@ 3.3 The role of the environment
Here, $\langle{\delta}(<R)\rangle$ and $\sigma_{\delta,\,R}(<R)$ represent the mean and standard deviation of the cosmic density contrast distribution, respectively. This formulation allows us to quantify how much a halo’s density contrast deviates from the sample mean: a positive $\delta_{\sigma,\,R}$ corresponds to an over-dense environment, whereas a negative value indicates an under-dense region.
Figure 7 shows the cumulative distribution function of the normalised density contrast for the bright and faint samples. The solid lines represent the median of the bootstrapped sample, while the shaded regions indicate the uncertainties. A striking difference emerges between the two populations, visible across all panels (from left to right, where $\delta_{\sigma,,R}$ is computed at different radii). The X-ray bright sample is preferentially located in over-dense environments, naturally explaining its enhanced ability to accrete gas over cosmic time. Additionally, we find a statistically significant preference for higher richness (i.e. the number of gravitationally bound subhalos) in the bright sample, a trend that persists at higher redshifts.
@@
Figure 9: Hot gas fraction as a function of the halo mass. The points are colour-coded for the residuals about the median of different properties as a function of the halo mass. The properties are: the BH mass $\Delta\log M_{\mathrm{BH}}$, the total metallicity $\Delta\log(Z_{500}/Z_{\odot})$, the total temperature $\Delta\log(k_{B}T_{500})$, and the richness $\Delta\log N_{\mathrm{gal}}$. In the bottom panels, we plot the Spearman coefficient between the hot gas fraction and the property in halo mass bins. The shaded band regions represent the bins where the p-value$<0.05$.
@@ 3.4.1 The SMBH properties
The first panel in Fig. 9 illustrates the trend of the $f_{\mathrm{gas}}-M_{\mathrm{halo}}$ relation when the central SMBH properties are considered. We colour-code the points with the residuals about the running median in halo mass bins of the BH mass $\Delta\log M_{\mathrm{BH}}$. We do not show the instantaneous BH accretion rate $\Delta\log\dot{M}_{\mathrm{BH}}$ or the number of BH–BH mergers $\Delta\log N_{\mathrm{(BH-BH)merger}}$, as they do not correlate. The strongest anti-correlations with the hot gas fraction are found for the BH mass ($\rho_{\{X,Y\}}\geq-0.5$). We interpret this in light of the different assembly histories that lead to different gas fractions at the same halo mass. Gas-poor halos tend to have massive central BHs. The only two mass-accretion channels for simulated BHs are either through multiple mergers (more than average) or a steady gas accretion, both returning thermal energy to their surroundings proportionally to the BH mass excess. However, the distribution of mergers does not support the former scenario. We argue that gas-poor halos have massive BHs because they depleted their gas content due to the AGN feeding/feedback mechanism and star formation, while no gas inflow has replenished its reservoir. However, gas-rich halos have on average lower BH masses since their rapid recent halo and gas accretion have not yet been followed by the BH activity (see also Fig. 6). Booth & Schaye (2010, 2011) found that the BH mass depends on the DM potential and evolution, rather than the properties of the stellar component. It is interesting to notice how very little the instantaneous BH accretion rate is affected by the hot gas fraction in the halo. To trigger this power, it is necessary to funnel the gas to the halo centre, as shown in Fig. 2.
It remains unclear to what extent the tight connection between central BH mass and hot gas fraction is a direct consequence of the simplified thermal (two-mode) AGN feedback model implemented in Magneticum (see Sect. 2.1 for details). This model prescribes a thermal energy injection into the surrounding gas particles, proportional to the mass accretion rate while neglecting other forms of feedback, such as non-isotropic outflows and jets. Similar trends are observed in other state-of-the-art cosmological simulations that employ comparable AGN feedback models, such as EAGLE and IllustrisTNG (Davies et al. 2020). A crucial step toward a more comprehensive understanding of these dependencies will be the inclusion of more sophisticated AGN feedback prescriptions in cosmological simulations (e.g. Sala et al. 2024; Rennehan et al. 2024).
@@ 3.4.2 The stellar properties
In the second and third panels at the top of Fig. 9, we show stellar properties, such as the stellar mass in the BCG with central halo $\Delta\log M_{\mathrm{\star,\,CEN}}$ and the mean stellar formation redshift $\Delta\log z_{\mathrm{\star,\,CEN}}$, and the fossilness $\Delta\log M_{1,2}$. These quantities are defined according to the star particles bound to the central subhalo in each group or cluster. The fossilness is defined as the logarithmic difference between the stellar masses of the BCG and the second most massive satellite in the halo. This definition often substitutes the magnitude gap used in observation to define the difference between the magnitudes of the central galaxy with the satellites. Such a definition shall describe how evolved the galaxy population is in a halo: a recently accreted or merged group will most likely have more massive orbiting satellites than a more relaxed system. We point out that this feature is vastly present in fossil/non-fossil groups (Jones et al. 2003). We define the BCG and its stellar halo as the subhalo 0 in the Subfind catalogue for each group. This corresponds to stars bound to the central galaxy and intracluster or intragroup light (generally, called ICL Montes 2022). Stars in the ICL are stars bound to the cluster potential, rather than a single galaxy. Other than yielding different dynamical properties to the BCG stars (e.g. Marini et al. 2021), the ICL is expected to have fallen into the cluster potential after secular stripping of the infalling satellite galaxies in the halo (Contini et al. 2014; Contini 2021; Brown et al. 2024; Savino et al. 2025). Hence, we expect larger stellar halos for systems with older formation times like the faint sample.
@@ 3.4.3 IGrM, ICM, and satellite properties
Metals are forged during the stellar evolution (in the cores of stars and SNe explosion) and the long-term effect of the AGN feedback re-distributes them across the halos (Biffi et al. 2017; Dolag et al. 2017). Therefore, older galaxy groups and clusters, with a larger stellar mass, will have a larger contribution from metals, compared to younger systems still assembling. Such contribution is significant only when the metallicity is measured up to larger radii since we can account for all the metals that have been moved outwards due to gas mixing. However, if the metallicity is measured in the core (see discussion in Marini et al. 2024) we observe no significant correlation with the X-ray luminosity (and, thus, gas fraction). Interestingly, the Spearman coefficient indicates that this effect increases with increasing mass. Similarly, within halos of constant mass, the non-zero Spearman coefficient for the temperature suggests the existence of a multiphase gas in gas-rich halos that has not yet been influenced by gravitational or non-gravitational heating processes. In older and X-ray fainter systems, the temperature is on average higher because the remaining gas has thermalised with the large potential well of the group. Conversely, measuring the number of satellites in $R_{500}$ indicates a higher fraction of galaxies surrounding gas-rich halos, and vice-versa (e.g. Puddu & Andreon 2022). This is consistent with a denser environment and recent mergers.
We point out that due to the substantially different availability of gas, the star formation rate of central galaxies is also expected to change significantly between (X-ray) bright and faint systems. Vladutescu-Zopp et al. (2025) investigated this effect for a higher resolution box than our own (i.e. Box4/uhr) finding star-forming central galaxies to be generally X-ray brighter than quiescent ones. Recent results from the synthetic eROSITA data (Oppenheimer et al. 2020; Vladutescu-Zopp et al. 2025) and eFEDS data also indicate the presence of such a correlation (Comparat et al. 2022; Chadayammuri et al. 2022), while the eRASS:4 stacking results contradict this picture (Zhang et al. 2025).

### v2

@@ 3.1 Build-up of the gas content
One might wonder how representative the hot gas fraction is of the total gas content in a group or cluster of galaxies. To address this, we analyse the expected hot gas fraction (hot gas mass over halo mass within $R_{500}$) as a function of the total gas fraction (hot and cold gas over halo mass within $R_{500}$) in Fig. 4. It is worth mentioning that Magneticum’s predictions on the $f_{\mathrm{gas}}-M_{halo}$ relation are in excellent agreement with the median and scatter of the hot gas fraction from eROSITA stacked groups (Popesso et al. 2024b) and the XXL groups and clusters with halo mass measurements from weak lensing (Akino et al. 2022). We divide the LC30 sample into logarithmic halo mass bins of $M_{500}$, showing their scatter (coloured shaded regions), and plot the 1:1 scaling relation in dashed black. If the relation follows the 1:1 line, this implies that all gas in the halo is hot. In the lateral panels, we plot the number counts of halos in the bins we used to calculate the median and error.
We deduce that if a halo has a low total gas fraction compared to haloes with a similar mass, most (if not all) of it is in the hot phase. The gravitational potential well compresses the hot plasma, increasing its kinetic energy, which is later converted into thermal energy. As a result, larger halos tend to heat more gas. In the clusters regime, this process converts most of the gas into a hot, X-ray-bright component. However, in less massive halos, the colder (star-forming) gas component can survive longer within $R_{500}$. Consequently, a larger total gas fraction does not necessarily imply a well-mixed hot halo. We argue that measuring only the X-ray-emitting gas in the group regime ($M_{500}<10^{13.5}M_{\odot}$) could lead to a significant underestimation of the total gas fraction (up to 50%) if this effect is not considered – at least in the $f_{\mathrm{gas,\,tot}}>0.05$ regime of our simulations. This result is robustly predicted by the number of halos that are estimated to have $f_{\mathrm{gas,\,tot}}$ in the regime $0.05-0.10$, which represent $\geq 50\%$ of the groups in $M_{500}<10^{13.5}M_{\odot}$ (see top histograms). In this regard, recent studies on the gas fraction have called for stronger AGN feedback than most hydrodynamical simulations (but not Magneticum, see Popesso et al. 2024c) based on evidence collected from the cosmic shear, X-ray, and SZ measurements (e.g. Preston et al. 2023; Amon et al. 2023; Bigwood et al. 2024; McCarthy et al. 2024; Hadzhiyska et al. 2024; Ferreira et al. 2024; La Posta et al. 2024). However, we highlight here the need to account for complementary techniques to include cold gas (e.g. measuring the MgII covering factor; Anand et al. 2021) in Milky Way-mass halos.
@@
Furthermore, the mass accretion history is strikingly different in the two samples (right panels in Fig. 5). We overplot the self-similar evolution of a DM halo in the given mass range at $z=0$ expected in the range of concentration $c_{500}$ [^4] The DM concentration $c_{\Delta}$ defines how centrally concentrated the DM is in a halo. Analytically it reads $c_{\Delta}=R_{\Delta}/R_{s}$ where $R_{s}$ is the scale radius in an NFW profile (Navarro et al. 1997). of our simulated halos $103<c_{500}<$ 10 (Diemer et al. 2013; Diemer 2018). Faint halos typically gain most of their mass early on, with later phases marked by gas depletion due to star formation, AGN activity, and stellar feedback. Interestingly, the halo mass growth experiences a slowdown in the faint sample at the epoch when the baryon fraction begins to decline. Such a decline can indeed be connected with a general failure in replenishing the gas content in the halo, which will reduce in the future due to various feedback and star formation. However, their overall evolution is consistent with the pseudo-evolution of a stationary NFW density profile (shown in Fig. 5), changing due to the evolution of a halo mass defined by a density threshold $\Delta=500$ that decreases with time. In contrast, bright halos exhibit steady exponential accretion at all times, removing the $1$ dex difference in halo mass originally present at $z=3.5$. A visual inspection of the bright halos at this epoch reveals that they undergo major mergers, whereas the faint ones exhibit very few substructures and display a regular, symmetric X-ray emission around the central galaxy –reminiscent of fossil groups (Aguerri & Zarattini 2021).
@@
Figure 7: Cumulative density distribution of the normalised density contrast for bright (orange) and faint (blue) halos at three redshifts. The solid lines are the median of the bootstrapped sample (shaded lines) for halos within $10^{13}-10^{13.5}M_{\odot}$ selected at $z=0$. From left to right, the normalised density contrast is calculated for different radii, namely 1, 2, and 5 cMpc/h. In the top-left corner of each panel, we also report its correspondence to the median $R_{500}$ in the sample. We overplot a vertical and horizontal line to guide the eye.
@@ 3.2 The role of the central SMBH
The activity of the central and satellite SMBHs can severely impact halos across their cosmic time. While AGN feedback heats and displaces the gas, altering the halo’s thermodynamic state (Eckert et al. 2021), many simulations (Steinborn et al. 2016, 2018; Castro et al. 2021) have shown it is insufficient to eject gas entirely from the halo at late times. In reality, most of the strong AGN activity episodes happen at the early phases of the halos assembly ($z=2-3$, see for example Castro et al. 2021; Ragagnin et al. 2022; Ayromlou et al. 2023) which can cause an efficient displacement of the gas well beyond the virial radius at the group scales. On the other hand, DM is less sensitive to the gas displacement (and completely insensitive to the thermal feedback), leading to a continuous accretion even at later times. Biffi et al. (2018) and Marini et al. (2024, 2025) showed that Magneticum correctly reproduces the AGN luminosity function at $z=0$ and in LC30, respectively. Therefore, we expect the simulated AGN population to be representative of the real one.
@@
The BH evolution demonstrates that the feedback activity of the faint sample steadily declines over time, with no recent episodes accounting for the decrease of the gas content around $z\simeq 0-1.5$. In contrast, the bright sample exhibits a rise in AGN activity at the epoch when the baryonic fraction changes ($z\simeq 1-1.5$). If a major merger occurs at this redshift, the instantaneous availability of gas in the halo fuels the central SMBH, triggering a thermal energy release in the bright sample. The BH accretion is more efficient and less turbulent during the accretion of cold gas particles, although both cold and hot gas phases contribute to the feeding process. Because cold gas is accreted more efficiently than hot gas, (i) it is quickly depleted and (ii) this selective removal of low-entropy gas boosts the overall temperature (McCarthy et al. 2011). In response, the SMBH releases thermal energy, which, while insufficient to fully deplete the gas reservoir of the halo, significantly increases the hot gas fraction.
@@ 3.3 The role of the environment
Here, $\langle{\delta}(<R)\rangle$ and $\sigma_{\delta,\,R}(<R)$ represent the mean and standard deviation of the cosmic density contrast distribution, respectively. This formulation quantifies how much a halo’s density contrast deviates from the sample mean: a positive $\delta_{\sigma,\,R}$ corresponds to an over-dense environment, whereas a negative value indicates an under-dense region.
Figure 7 shows the cumulative distribution function of the normalised density contrast for the bright and faint samples. The solid lines represent the median of the bootstrapped sample, while the shaded regions indicate the uncertainties. A striking difference emerges between the two populations, visible across all panels (from left to right, where $\delta_{\sigma,R}$ is computed at different radii). The X-ray bright sample is preferentially located in median-dense/over-dense environments, naturally explaining its enhanced ability to accrete gas over cosmic time. Conversely, the X-ray faint sample is consistently in the under-dense region of the diagrams at all $z$. Additionally, we find a statistically significant preference for higher richness (i.e. the number of gravitationally bound subhalos) in the bright sample, a trend that persists at higher redshifts.
@@
Figure 9: Hot gas fraction as a function of the halo mass. The points are colour-coded for the residuals about the median of different properties as a function of the halo mass. The properties are: the BH mass $\Delta\log M_{\mathrm{BH}}$, the BCG stellar mass $\Delta\log M_{\mathrm{\star,CEN}}$, the mean stellar formation redshift $\Delta z_{\mathrm{\star,CEN}}$, the magnitude gap $\Delta\log M_{\mathrm{1,2}}$, the richness $\Delta\log N_{\mathrm{SAT}}$, the total metallicity $\Delta\log(Z_{500}/Z_{\odot})$, and the total temperature $\Delta\log(k_{B}T_{500})$. In the bottom panels, we plot the Spearman coefficient between the hot gas fraction and the property in halo mass bins. The shaded band regions represent the bins where the p-value$<0.05$. The analogue plot with X-ray luminosity in the soft band is in Fig. 11.
@@ 3.4.1 The SMBH properties
The first panel in Fig. 9 illustrates the trend of the $f_{\mathrm{gas}}-M_{\mathrm{halo}}$ relation when the central SMBH properties are considered. We colour-code the points with the residuals about the running median in halo mass bins of the BH mass $\Delta\log M_{\mathrm{BH}}$. We do not show the instantaneous BH accretion rate $\Delta\log\dot{M}_{\mathrm{BH}}$ or the number of BH–BH mergers $\Delta\log N_{\mathrm{(BH-BH)merger}}$, as they do not correlate. The strongest anti-correlations with the hot gas fraction are found for the BH mass ($\rho_{\{X,Y\}}\leq-0.5$). We interpret this in light of the different assembly histories that lead to different gas fractions at the same halo mass. Gas-poor halos tend to have massive central BHs. The only two mass-accretion channels for simulated BHs are either through multiple mergers (more than average) or a steady gas accretion, both returning thermal energy to their surroundings proportionally to the BH mass excess. However, the distribution of mergers does not support the former scenario. We argue that gas-poor halos have massive BHs because they depleted their gas content due to the AGN feeding/feedback mechanism and star formation, while no gas inflow has replenished their reservoir. However, gas-rich halos have on average lower BH masses since their rapid recent halo and gas accretion have not yet been followed by the BH activity (see also Fig. 6). Booth & Schaye (2010, 2011) found that the BH mass depends on the DM potential and evolution, rather than the properties of the stellar component. It is interesting to notice how very little the instantaneous BH accretion rate is affected by the hot gas fraction in the halo. To trigger this power, it is necessary to funnel the gas to the halo centre, as shown in Fig. 2.
It remains unclear to what extent the tight connection between central BH mass and hot gas fraction is a direct consequence of the simplified thermal (two-mode) AGN feedback model implemented in Magneticum (see Sect. 2.1 for details). This model prescribes a thermal energy injection into the surrounding gas particles, proportional to the mass accretion rate, while neglecting other forms of feedback, such as non-isotropic outflows and jets. Similar trends are observed in other state-of-the-art cosmological simulations that employ comparable AGN feedback models, such as EAGLE and IllustrisTNG (Davies et al. 2020). A crucial step toward a more comprehensive understanding of these dependencies will be the inclusion of more sophisticated AGN feedback prescriptions in cosmological simulations (e.g. Sala et al. 2024; Rennehan et al. 2024).
@@ 3.4.2 The stellar properties
In the second and third panels at the top of Fig. 9, we show stellar properties, such as the stellar mass in the BCG with central halo $\Delta\log M_{\mathrm{\star,\,CEN}}$ and the mean stellar formation redshift $\Delta\log z_{\mathrm{\star,\,CEN}}$, and the fossilness $\Delta\log M_{1,2}$. These quantities are defined according to the star particles bound to the central subhalo in each group or cluster. The fossilness is defined as the logarithmic difference between the stellar masses of the BCG and the second most massive satellite in the halo. This definition often substitutes the magnitude gap used in observation to define the difference between the magnitudes of the central galaxy and the satellites. Such a definition shall describe how evolved the galaxy population is in a halo: a recently accreted or merged group has most likely more massive orbiting satellites than a more relaxed system. We point out that this feature is vastly present in fossil/non-fossil groups (Jones et al. 2003). We define the BCG and its stellar halo as the subhalo 0 in the Subfind catalogue for each group. This corresponds to stars bound to the central galaxy and intracluster or intragroup light (generally, called ICL Montes 2022). Stars in the ICL are stars bound to the cluster potential, rather than a single galaxy. Other than yielding different dynamical properties to the BCG stars (e.g. Marini et al. 2021), the ICL is expected to have fallen into the cluster potential after secular stripping of the infalling satellite galaxies in the halo (Contini et al. 2014; Contini 2021; Brown et al. 2024; Savino et al. 2025). Hence, we expect larger stellar halos for systems with older formation times, like the faint sample.
@@ 3.4.3 IGrM, ICM, and satellite properties
Metals are forged during the stellar evolution (in the cores of stars and SNe explosions) and the long-term effect of the AGN feedback re-distributes them across the halos (Biffi et al. 2017; Dolag et al. 2017). Therefore, older galaxy groups and clusters, with a larger stellar mass, have a larger contribution from metals, compared to younger systems still assembling. Such a contribution is significant only when the metallicity is measured up to larger radii, since we can account for all the metals that have been moved outwards due to gas mixing. However, if the metallicity is measured in the core (see discussion in Marini et al. 2024) we observe no significant correlation with the X-ray luminosity (and, thus, gas fraction). Interestingly, the Spearman coefficient indicates that this effect increases with increasing mass. Similarly, within halos of constant mass, the non-zero Spearman coefficient for the temperature suggests the existence of a multiphase gas in gas-rich halos that has not yet been influenced by gravitational or non-gravitational heating processes. In older and X-ray fainter systems, the temperature is on average higher because the remaining gas has thermalised with the large potential well of the group. Conversely, measuring the number of satellites in $R_{500}$ indicates a higher fraction of galaxies surrounding gas-rich halos, and vice-versa (e.g. Puddu & Andreon 2022). This is consistent with a denser environment and recent mergers.
We point out that due to the substantially different availability of gas, the star formation rate of central galaxies is also expected to change significantly between (X-ray) bright and faint systems. Vladutescu-Zopp et al. (2025) investigated this effect for a higher resolution box than our own (i.e. Box4/hr), finding star-forming central galaxies to be generally X-ray brighter than quiescent ones. Recent results from the synthetic eROSITA data (Oppenheimer et al. 2020; Vladutescu-Zopp et al. 2025) and eFEDS data also indicate the presence of such a correlation (Comparat et al. 2022; Chadayammuri et al. 2022), while the eRASS:4 stacking results contradict this picture (Zhang et al. 2025).